# The influence of the 3D Galactic gas structure on cosmic-ray transport and -ray emission

### v1

@@ Abstract
Context. Cosmic rays (CRs) play a major role in the dynamics of the interstellar medium (ISM). Their interactions and transport ionize, heat, and push the ISM thereby coupling different regions of it. The spatial distribution of CRs depends on the distribution of their sources as well as the ISM constituents they interact with, such as gas, starlight, and magnetic fields. Particularly, gas interacts closely with CRs, influencing CR fluxes and -ray emission.
Aims. We illustrate the influence of 3D gas structures on CR transport and -ray emission.
Methods. We use the PICARD code and multiple samples of recent 3D reconstructions of the HI and H2 Galactic gas constituents to investigate the impact on the transport of CRs and emission of -rays.
Results. We find the necessary transport parameters to reproduce local measurements of CR fluxes, and see that they depend on the local distribution of gas density and structure. The distribution of CR fluxes exhibits energy-dependent structures that vary for all CR species due to their corresponding loss processes. Regions of enhanced secondary (primary) species are spatially correlated (anti-correlated) with the gas density. We observe a high sensitivity of the -ray emission on the contrast of gas structures, as those determine the 3D spatial distributions of hadronic interactions and bremsstrahlung. We find that corresponding gas-induced structures in the distribution of CR electrons are also visible in Inverse Compton (IC) emission. Due to the aforementioned sensitivity, the analysis of CR data for CR sources and transport parameters requires the usage of accurate 3D gas maps.
Key Words.:
Diffusion ā Methods: numerical ā cosmic rays ā ISM: kinematics and dynamics ā Gamma rays: ISM1 Introduction
@@
Coscmic rays (CRs) are an important component of the Galactic interstellar medium (ISM). They interact with other constituents of the ISM such as gas, dust, turbulent magnetic fields and interstellar radiation. Furthermore, CRs are believed to affect the chemical and dynamical evolution of the ISM from small (star-forming) to large, Galactic scales (Padovani etĀ al., 2020; Gabici, 2022; Phan etĀ al., 2023; Simpson etĀ al., 2023). Several studies suggest that CRs drive Galactic winds and, thus, could play an essential role in the evolution history of the Milky Way (MW) (Farcy etĀ al., 2022; Girichidis etĀ al., 2022).
@@
Despite the importance of CRs in the Galactic ecosystem, the transport of these particles is still far from being fully understood. Analyses of CR secondary-to-primary ratios (e.g. B/C and 10Be/9Be) indicate that these particles propagate diffusively within a magnetized halo (with a typical height of around 3-6ākpc) surrounding the Galactic disk (Evoli etĀ al., 2020; Maurin etĀ al., 2022; Jacobs etĀ al., 2023). The diffusion of CRs is thought to be due to their interactions with magnetic turbulence (the turbulent component of the Galactic magnetic field). In this respect, Galactic CR modelling formally follows from coupling the transport of CRs to magnetic turbulence (Schlickeiser, 2002; Blasi, 2013), using a set of non-linear coupled differential equations. Solving these coupled equations remains challenging in many different ways. For example, it is not yet clear whether magnetic turbulence is predominantly generated from the cascade of large-scale turbulence injected by supernova explosions or from CRs themselves via the streaming instability (Skilling, 1975; Nava etĀ al., 2016; Jacobs etĀ al., 2022). Several recent analyses, in fact, indicate that both mechanisms are important (Thomas & Pfrommer, 2019; Thomas etĀ al., 2020) and in different energy ranges they lead to features in the CR spectra (GĆ©nolini etĀ al., 2017; Evoli etĀ al., 2018). Another difficulty in modelling Galactic CR transport and magnetic turbulence self-consistently is our lack of knowledge on the large-scale magnetic field of the MW (Beck, 2015). More precise 3D models of the Galactic magnetic field (Jansson & Farrar, 2012a, b) would be essential for a better understanding of CR propagation (Haverkorn etĀ al., 2019) since the diffusion coefficients of CRs parallel and perpendicular to the field lines are expected to differ, making the Galactic distribution of CRs sensitive to the exact geometry of the large-scale magnetic field (Cerri etĀ al., 2017; Mertsch, 2020; Giacinti & Semikoz, 2023).
@@
Given all the above-mentioned complexities, the propagation of CRs is normally modelled in a linear fashion, via the CR transport equation assuming parametric forms of the diffusion coefficients and other transport and acceleration parameters such as the advection speed (due to plasma waves or Galactic winds) or the diffusion coefficient in momentum space (for the process of re-acceleration). In fact, the growing amount of high-precision CR data has led to the development and refinement of multiple sophisticated CR transport and interaction codes such as GALPROP (Porter etĀ al., 2022), DRAGON (Evoli etĀ al., 2017), USINE (Maurin, 2020) and PICARD (Kissmann, 2014). These codes provide numerical or semi-analytical solutions for the CR transport equation, taking into account many different physical processes not only for transport but also for particle acceleration and interactions between CRs and the ISM (resulting in energy loss and production of secondary CRs via spallation). The combination of these transport codes and high-precision CR data (Adriani etĀ al., 2014, 2022; Aguilar etĀ al., 2023), covering energy ranges from the MeV to the TeV scale and beyond, has been extensively used to extract corresponding transport and acceleration parameters (Strong etĀ al., 2007a; Grenier etĀ al., 2015; Gabici etĀ al., 2019).
@@
In many of these CR transport codes, gas distributions are, however, modelled through the introduction of axisymmetric Galacto-centric models. Since gas serves as a target for spallation reactions of primary CRs producing secondary CRs, approximate 2D gas distributions and more realistic 3D ones are expected to yield different estimates for secondary-to-primary ratios for otherwise identical transport parameters. Conversely, different estimates of the transport parameters are expected between 2D and 3D gas models when inferring them from data on CR secondary-to-primary ratios. This is observed when a more structured 3D model is used for the inference of transport parameters instead of a smooth axisymmetric model, yielding a lower diffusion coefficient and a reduced Alfvèn speed (see e.g., Jóhannesson et al. (2018), but also Table 1 here).
@@
Furthermore, interactions between CRs and gas can result in observable diffuse -ray emissions, i.e.Ā via decays of (produced in hadronic interactions of CRs with interstellar gas nuclei) or via bremsstrahlung by CR electrons. Diffuse -ray data can, therefore, be used to derive the Galactic distribution of CRs assuming models for the 3D gas distributions (Ackermann etĀ al., 2012; Tibaldo etĀ al., 2021). Despite their relevance, 3D gas distributions are still not commonly used in studies of Galactic CR transport.
@@
One of the main methods for deriving 3D gas distributions is to analyze data of spectral emission and absorption lines. Atomic hydrogen (HI), for instance, can be traced through the 21-cm emission line. Since gas at different locations along a line of sight in general has different relative velocities with respect to the observer due to Galactic rotation, it emits this line at different Doppler-shifted frequencies. This means that, given a model for Galactic rotation, data of this emission line can be used to reconstruct a model for the 3D distribution of HI. A similar approach can be applied to the emission line of 12CO to trace molecular hydrogen (H2) assuming a conversion factor between 12CO and H2 gas density. However, given our vantage point in the Galaxy, many of these reconstructed gas maps suffer from a number of ambiguities. For instance, the peculiar motion of gas (i.e.Ā random motions of gas, on top of the large-scale rotational gas flow caused e.g. by stellar winds and supernova explosions) introduce artificial structures stretching along the lines of sight into the 3D gas reconstructions, known as the finger-of-god effect (see e.g., Nakanishi & Sofue, 2003, 2006; Pohl etĀ al., 2008). Recently it has been suggested that this problem can be cured by taking the spatial correlations of gas into account (Mertsch, P. & Vittino, A., 2021; Mertsch, P. & Phan, V. H. M., 2023). Incorporating spatial correlations into 3D gas reconstructions turns those into high-dimensional Bayesian inference problems. Such have been addressed via information field theory (EnĆlin etĀ al., 2009; EnĆlin, 2019) in Mertsch, P. & Vittino, A. (2021) and Mertsch, P. & Phan, V. H. M. (2023) by using the NIFTy package111github.com/NIFTy-PPL/NIFTy(Selig etĀ al., 2013; Steininger etĀ al., 2019; Arras etĀ al., 2019; Edenhofer etĀ al., 2024). This has not only provided improved 3D gas reconstructions but also quantified their remaining uncertainties.
@@
In this work, we aim to study how the 3D gas structures, as seen in such improved 3D gas reconstructions, influence CR transport and -ray emission. We model this by introducing 3D reconstructed maps into diffusive CR transport simulations. For comparison, we also produce an axisymmetric version of the same 3D reconstructed maps. We use both models to infer transport parameters, and to simulate the energy-dependent spatial distribution of CRs and of -ray emission with the PICARD code.
@@ 2 Methodology
aā Normalisation flux given in units of [m-2s-1sr-1(GeV/nucleon)-1] |
@@
In the PICARD code, a steady state solution to the CR transport equation for each nucleus is obtained on a numerical grid either using axisymmetric cylindrical coordinates or a three-dimensional Cartesian grid in space ā as described for our models below ā and one energy dimension, where we use a logarithmically equidistant axis, ranging from 10āMeV to 1āPeV in 127 grid points. Our setup is based on previous axisymmetric models from Ackermann etĀ al. (2012), i.e. their model Y4201505. In particular, we are using the axisymmetric cosmic-ray source distribution by Yusifov & Küçük (2004), which we truncated beyond a Galactocentric radius of 15ākpc as also discussed in Kissmann etĀ al. (2015). The transport parameters are the same as in Y4201505 apart from AlfvĆ©n speed and normalisation and energy dependence of spatial diffusion, as discussed below. Additionally, the models for the magnetic field and the radiation field are also axisymmetric as also used in Ackermann etĀ al. (2012). Correspondingly, the new three-dimensional gas reconstructions are the only non-axisymmetric components considered in the transport equation. We use the same set of spallation cross-sections and energy loss process as GALPROP v54222galprop.stanford.edu/download/manuals/galprop_v54.pdf (see e.g., Strong & Moskalenko, 1998; Moskalenko & Strong, 1998; Moskalenko etĀ al., 2002).
@@
Nuclei are coupled through spallation reactions which lead to corresponding loss and source terms. For the handling of the related nuclear network see Kissmann etĀ al. (2015). Thus, for particles resulting from interactions with Galactic gas, defines the interaction regions producing secondaries.
@@
We adapt the recently reconstructed 3D maps of molecular (H2) (Mertsch, P. & Vittino, A., 2021) and atomic (HI) (Mertsch, P. & Phan, V. H. M., 2023) hydrogen into the PICARD simulations. These reconstructions were performed following two different gas flow models, namely BEG03 (based on the results of a smoothed particle hydrodynamics simulation) and SBM15 (based on a semi-analytical model for gas-carrying orbits) as introduced in the original articles. Since we are interested in the distribution of 3D structures, both are equally good candidates for the purpose of this work. Here, we selected the HI and H2 maps based on the BEG03 gas flow model. These maps come in the form of 32 H2 and 16 HI independent posterior sample maps. Out of the 512 possible combinations of H2 and HI sample pairs, we subsample by randomly choosing 20 map pairs to account for fluctuations in the gas structure. To load each pair of gas maps into the simulations, we interpolate them from their original reconstruction grid onto the PICARD simulation grid. The gas maps are then added with the analytical model of ionised (HII) gas used in PICARD to generate the total distribution of protons in the galaxy . This is calculated as
@@
From these maps we produce two simulation setups. For the first one, from hereon referred to as the 2D model, we average over the selected sample maps to produce a reconstruction mean map. We interpolate this map from the reconstruction grid to a cylindrical grid . To produce an axisymmetric model, we average the result over the azimuth . The resulting 2D distribution is given on a grid with 128 grid points in the dimension and 65 grid points in the dimension. The grid spans from 0ākpc to 20ākpc in and from -4ākpc to 4ākpc in . For the second setup, from hereon referred to as the 3D model, we interpolate the 20 samples directly onto the Cartesian simulation grid . This grid is defined by 257 grid points in both and dimensions spanning the same range from -20ākpc to 20ākpc. The axis is identical to the one in the 2D model. As a result, both configurations possess the same resolution.
@@
Both setups are illustrated in Fig 1 where we show the resulting surface mass density () of the reconstruction mean map and the uncertainty (e.g one standard deviation) computed over the samples, as well as the surface mass density of the 2D model map.
@@
Different gas models cause the simulated secondary to primary ratios at Earth to differ when compared to data (i.e. B/C ratio data). To understand these differences, we heuristically calibrate our transport parameters by modifying those related to the strength and speed of CR diffusion. Namely, the diffusion coefficient , its corresponding indices and , and the AlfvĆØn speed . We perform this calibration by running multiple simulations with slightly different values for the parameters thereof, until finding a rough by-eye agreement between the simulated B/C ratios and AMS-02 data (Aguilar etĀ al., 2018). We do this for both the 2D and 3D models using the axisymmetric and reconstruction mean maps shown in Fig. 1, respectively. In this work, the heaviest element considered in the simulations is Silicon, where contributions from heavier elements are neglected, as they impact the distribution of Carbon and Boron isotopes by less than a percent. The resulting set of transport parameters are presented in Table 1.
Nuclei are coupled through spallation reactions which lead to corresponding loss and source terms. For the handling of the related nuclear network see Kissmann etĀ al. (2015). Thus, for particles resulting from interactions with Galactic gas, defines the interaction regions producing secondaries.
@@
Diffuse -ray emission is computed considering the contributions from -decay, bremsstrahlung, and IC scattering. For this calculation, transport results for relevant CR species (hydrogen, helium, electrons and positrons) are used with the corresponding cross-sections to generate emissivities at the specified energies of interest (see e.g, Strong & Moskalenko, 1998; Strong etĀ al., 2000). For -decay and bremsstrahlung, specific emissivities are produced which are later multiplied with the total proton distribution to obtain the final -ray emissivities in the Galaxy. For these simulations, the spatial grid is the same as in the transport step and CR energies range from 10āGeV to 1āTeV in 10 logarithmically equidistant grid points. After the diffuse emission is computed, the specific emissivities for -decay and bremsstrahlung are interpolated onto a Cartesian grid with the same spatial resolution as the sample maps, to multiply them with the total gas distribution. Such grid preserves the full resolution of the sample maps, spanning the same dimensions as that of the 3D model, with (640,640,128) grid points in the () dimensions. From the local emission, all-sky -ray emission follows via line-of-sight (LOS) integration.

### v2

@@ Abstract
Cosmic rays (CRs) play a major role in the dynamics of the interstellar medium (ISM). Their interactions and transport ionize, heat, and push the ISM thereby coupling different regions of it. The spatial distribution of CRs depends on the distribution of their sources as well as the ISM constituents they interact with, such as gas, starlight, and magnetic fields. Particularly, gas influences CR fluxes and -ray emission. We illustrate the influence of realistic and largely structured 3D gas distributions on CR transport and -ray emission, by studying their correlation using the PICARD code and multiple samples of recent 3D reconstructions of the HI and H2 Galactic gas constituents. We adjust the diffusion coefficient and Alfvén speed to reproduce local measurements of B/C abundances and find that these parameters depend non-linearly on the local distribution of gas. When simulating CR transport, the distributions of CR fluxes exhibit energy-dependent structures that vary for all CR species due to their corresponding loss processes. Regions of enhanced secondary (primary) species are spatially correlated (anti-correlated) with the gas density. Furthermore, we show that the morphology of gas clouds alone impacts CR flux predictions. For -ray emission, we observe a high sensitivity of the -ray emissivities to gas structures, as these determine the spatial distributions of hadronic interactions and bremsstrahlung. This way, we have for the first time calculated how well-defined uncertainties in a structured gas model propagate to CR transport and -ray emission.
keywords:
Diffusion – Cosmic rays – Gamma rays – Galactic gas – Numerical methodsOmm
[…]
1 Introduction
@@
Cosmic rays (CRs) are an important component of the Galactic interstellar medium (ISM). They interact with other constituents of the ISM such as gas, dust, turbulent magnetic fields and interstellar radiation. Furthermore, CRs are believed to affect the chemical and dynamical evolution of the ISM from small (star-forming) to large, Galactic scales (Padovani et al., 2020; Gabici, 2022; Phan et al., 2023; Simpson et al., 2023). Several studies suggest that CRs drive Galactic winds and, thus, could play an essential role in the evolution history of the Milky Way (MW) (Farcy et al., 2022; Girichidis et al., 2022).
@@
Despite the importance of CRs in the Galactic ecosystem, the transport of these particles is still far from being fully understood. Analyses of CR secondary-to-primary ratios (e.g. B/C and 10Be/9Be) indicate that these particles propagate diffusively within a magnetized halo (with a typical height of around 3-6 kpc) surrounding the Galactic disk (Evoli et al., 2020; Maurin et al., 2022; Jacobs et al., 2023). The diffusion of CRs is thought to be due to their interactions with magnetic turbulence (the turbulent component of the Galactic magnetic field). In this respect, Galactic CR modelling formally follows from coupling the transport of CRs to magnetic turbulence (Schlickeiser, 2002; Blasi, 2013), using a set of non-linear coupled differential equations. Solving these coupled equations remains challenging in many different ways. For example, it is not yet clear whether magnetic turbulence is predominantly generated from the cascade of large-scale turbulence injected by supernova explosions or from CRs themselves via the streaming instability (Skilling, 1975; Nava et al., 2016; Jacobs et al., 2022). Several recent analyses, in fact, indicate that both mechanisms are important (Thomas and Pfrommer, 2019; Thomas et al., 2020) and in different energy ranges they lead to features in the CR spectra (Génolini et al., 2017; Evoli et al., 2018). Another difficulty in modelling Galactic CR transport and magnetic turbulence self-consistently is our lack of knowledge on the large-scale magnetic field of the MW (Beck, 2015). More precise 3D models of the Galactic magnetic field (Jansson and Farrar, 2012a, b) would be essential for a better understanding of CR propagation (Haverkorn et al., 2019) since the diffusion coefficients of CRs parallel and perpendicular to the field lines are expected to differ, making the Galactic distribution of CRs sensitive to the exact geometry of the large-scale magnetic field (Cerri et al., 2017; Mertsch, 2020; Giacinti and Semikoz, 2023).
@@
Given all the above-mentioned complexities, the propagation of CRs is normally modelled in a linear fashion, via the CR transport equation assuming parametric forms of the diffusion coefficients and other transport and acceleration parameters such as the advection speed (due to plasma waves or Galactic winds) or the diffusion coefficient in momentum space (for the process of re-acceleration). In fact, the growing amount of high-precision CR data has led to the development and refinement of multiple sophisticated CR transport and interaction codes such as GALPROP (Porter et al., 2022), DRAGON (Evoli et al., 2017), USINE (Maurin, 2020) and PICARD (Kissmann, 2014). These codes provide numerical or semi-analytical solutions for the CR transport equation, taking into account many different physical processes not only for transport but also for particle acceleration and interactions between CRs and the ISM (resulting in energy loss and production of secondary CRs via spallation). The combination of these transport codes and high-precision CR data (Adriani et al., 2014, 2022; Aguilar et al., 2023), covering energy ranges from the MeV to the TeV scale and beyond, has been extensively used to extract corresponding transport and acceleration parameters (Strong et al., 2007; Grenier et al., 2015; Gabici et al., 2019).
@@
Many models using such codes use axisymmetric Galacto-centric gas distributions to simulate CR transport (see e.g., Evoli et al., 2012; Gaggero et al., 2015; Orlando, 2019; Widmark et al., 2023). Since gas serves as a target for spallation reactions of primary CRs producing secondary CRs, approximate 2D gas distributions and more realistic 3D ones are expected to yield different estimates for secondary-to-primary ratios for otherwise identical transport parameters. Conversely, different estimates of the transport parameters are expected between 2D and 3D gas models when inferring them from data on CR secondary-to-primary ratios. This is observed when a more structured 3D model is used for the inference of transport parameters instead of a smooth axisymmetric model, yielding a lower diffusion coefficient and a reduced Alfvén speed (see e.g., Jóhannesson et al. (2018), but also Table 1 here).
@@
Furthermore, interactions between CRs and gas can result in observable diffuse -ray emissions, i.e. via decays of (produced in hadronic interactions of CRs with interstellar gas nuclei) or via bremsstrahlung by CR electrons. Diffuse -ray data can, therefore, be used to derive the Galactic distribution of CRs assuming models for the 3D gas distributions (Ackermann et al., 2012; Tibaldo et al., 2021). Despite their relevance, 3D gas distributions are still not commonly used in studies of Galactic CR transport.
@@
One of the main methods for deriving 3D gas distributions is to analyze data of spectral emission and absorption lines. Atomic hydrogen (HI), for instance, can be traced through the 21-cm emission line. Since gas at different locations along a line of sight in general has different relative velocities with respect to the observer due to Galactic rotation, it emits this line at different Doppler-shifted frequencies. This means that, given a model for Galactic rotation, data of this emission line can be used to reconstruct a model for the 3D distribution of HI. A similar approach can be applied to the emission line of 12CO to trace molecular hydrogen (H2) assuming a conversion factor between 12CO and H2 gas density. However, given our vantage point in the Galaxy, many of these reconstructed gas maps suffer from a number of ambiguities. For instance, the peculiar motion of gas (i.e. random motions of gas, on top of the large-scale rotational gas flow caused e.g. by stellar winds and supernova explosions) introduce artificial structures stretching along the lines of sight into the 3D gas reconstructions, known as the finger-of-god effect (see e.g., Nakanishi and Sofue, 2003, 2006; Pohl et al., 2008). Recently it has been suggested that this problem can be cured by taking the spatial correlations of gas into account (Mertsch, P. and Vittino, A., 2021; Mertsch, P. and Phan, V. H. M., 2023). Incorporating spatial correlations into 3D gas reconstructions turns those into high-dimensional Bayesian inference problems. Such have been addressed via information field theory (Enßlin et al., 2009; Enßlin, 2019) in Mertsch, P. and Vittino, A. (2021) and Mertsch, P. and Phan, V. H. M. (2023) by using the NIFTy package111github.com/NIFTy-PPL/NIFTy(Selig et al., 2013; Steininger et al., 2019; Arras et al., 2019; Edenhofer et al., 2024). This has not only provided improved 3D gas reconstructions but also quantified their remaining uncertainties. An important feature of these gas reconstructions, when compared to analytical 3D models, is their higher dynamic range (i.e. local densities vary in a wider range) which leads to a higher contrast between over- and under-density regions.
@@
In this work, we study how 3D gas structures, as seen in such improved 3D gas reconstructions, influence CR transport and -ray emission. We model this by introducing the 3D reconstructed maps and an axisymmetric averaged version of the same 3D maps into CR transport simulations. We use both models to correlate gas densities to transport parameters (i.e. Diffusion coefficient and Alfvén speed), and to simulate the energy-dependent spatial distribution of CRs and of -ray emission with the PICARD code. In contrast to previous works (e.g. Jóhannesson et al., 2018), rather than aiming for a best-fit model for each investigated setup, we focus on a model-to-model comparison to quantify the impact of localized gas structures on the simulated distributions of CR fluxes and -ray emissivities. In particular for the latter, we ensure the same 3D gas distribution is used to perform the line-of-sight integration leading to all-sky -ray emission maps.
@@ 2 Methodology
| a Normalisation flux given in units of [m-2s-1sr-1(GeV/nucleon)-1] | ||
@@
In the PICARD code, a steady state solution to the CR transport equation for each nucleus is obtained on a numerical grid by either using an axisymmetric cylindrical coordinates or a three-dimensional Cartesian grid in space – as described for our models below – and one energy dimension, where we use a logarithmically equidistant axis, ranging from 10 MeV to 1 PeV in 127 grid points.
Our setup is based on a previous axisymmetric model from Ackermann et al. (2012), i.e. using the short-hand notation introduced therein, the model Y4201505. This designator refers to a halo size of kpc, radial boundaries at kpc, spin temperature K, and magnitude cut of . In particular, we use the axisymmetric cosmic-ray source distribution by Yusifov and Küçük (2004), which we truncate beyond a Galactocentric radius of 15 kpc as also discussed in Kissmann et al. (2015). The transport parameters are the same as in Y4201505 apart from the Alfvén speed, the energy dependence of spatial diffusion, and the particle normalizations, as discussed below. Additionally, the models for the magnetic field and the radiation field are also axisymmetric as also used in Ackermann et al. (2012). Correspondingly, the new three-dimensional gas reconstructions are the only non-axisymmetric components considered in the transport equation. We use the same set of spallation cross-sections and energy loss process as GALPROP v54222galprop.stanford.edu/download/manuals/galprop_v54.pdf (see e.g., Strong and Moskalenko, 1998; Moskalenko and Strong, 1998; Moskalenko et al., 2002).
@@
Nuclei are coupled through spallation reactions which lead to corresponding loss and source terms. For the handling of the related nuclear network see Kissmann et al. (2015). Thus, for particles resulting from interactions with Galactic gas, defines the interaction regions producing secondaries.
@@
We adapt the recently reconstructed 3D maps of molecular (H2) – based on the reconstruction of Galactic CO with a fixed cm-2(K km s-1)-1 – (Mertsch, P. and Vittino, A., 2021) and atomic (HI) (Mertsch, P. and Phan, V. H. M., 2023) hydrogen into the PICARD simulations. These reconstructions were performed following two different gas flow models, namely BEG03 (based on the results of a smoothed particle hydrodynamics simulation) and SBM15 (based on a semi-analytical model for gas-carrying orbits) as introduced in the original articles. Since we are interested in the distribution of 3D structures, both are equally good candidates for the purpose of this work. Here, we selected the HI and H2 maps based on the BEG03 gas flow model. These maps come in the form of 32 H2 and 16 HI independent posterior sample maps. Out of the 512 possible combinations of H2 and HI sample pairs, we subsample by randomly choosing 20 map pairs to account for fluctuations in the gas structure. To load each pair of gas maps into the simulations, we interpolate them from their original reconstruction grid onto the PICARD simulation grid. The gas maps are then added with the analytical model of ionised (HII) gas used in PICARD to generate the total distribution of protons in the galaxy . This is calculated as
@@
We assume that the distribution of ISM He follows the one of ISM H, i.e. we compute the corresponding distribution assuming the number density of He to be a fraction of 11% of . From these maps we produce two models. For the first one, from hereon referred to as the 2D model, we average over the selected sample maps to produce a reconstruction mean map. We interpolate this map from the reconstruction grid to a cylindrical grid with the same resolution in as that of and in the original reconstruction grid, as well as the same resolution in , and 720 grid points defining the axis. To produce an axisymmetric 2D model, we average the result over the azimuth and store the averaged value at the corresponding grid point for every . The resulting distribution is then interpolated to a 3D Cartesian simulation grid. For the second setup, from hereon referred to as the 3D model, we interpolate the 20 samples directly onto the same Cartesian simulation grid . The Cartesian simulation grid spans from -20 kpc to 20 kpc with 257 grid points in both and dimensions, and from -4 kpc to 4 kpc with 65 grid points in .
Given the multiple samples we study here, higher resolution setups become prohibitively expensive. To test the impact of the chosen spatial resolution, we investigate the differences between predictions using our default grid and a finer one, using twice the spatial-resolution (i.e. 513 grid points in and 129 grid points in ). We compare the simulated distributions of primary and secondary protons using both grids at the two main locations of interest for this work: Earth and the Galactic center. On the coarse grid, we find that the fluxes are on average higher by 2.5% without any changes to the spectral slope. Deviations are larger only at very localized gas clouds, raising to 10%, due to the shape of these clouds blurring out in the lower resolution. However, structures are not principally different. Thus the coarse grid proves sufficient for the purpose of this work.
@@
Both models are illustrated in Fig 1 where we show the resulting surface mass density () of the reconstruction mean map and the uncertainty (e.g one standard deviation) computed over the samples, as well as the surface mass density of the 2D model map.
@@
Different gas models cause the simulated secondary to primary ratios at Earth to differ when compared to data (i.e. B/C ratio data). To understand these differences, we heuristically calibrate our transport parameters by modifying those related to the strength and speed of CR diffusion. Namely, the diffusion coefficient , its corresponding indices and , and the Alfvén speed . We refrain from fitting many transport parameters for each model, because we are mainly interested on the impact of the gas only. We perform this calibration by running multiple simulations with slightly different values for the parameters thereof, until finding a rough by-eye agreement between the simulated B/C ratios and AMS-02 data (Aguilar et al., 2018). We do this for both the 2D and 3D models using the axisymmetric and reconstruction mean maps shown in Fig. 1, respectively. In this work, the heaviest element considered in the simulations is Silicon, where contributions from heavier elements are neglected, as they impact the distribution of Carbon and Boron isotopes by less than a percent. The resulting set of transport parameters are presented in Table 1.
@@
Diffuse -ray emission is computed considering the contributions from -decay, bremsstrahlung, and IC scattering. For this calculation, transport results for relevant CR species (hydrogen, helium, electrons and positrons) are used with the corresponding cross-sections to generate emissivities at the specified energies of interest (see e.g, Strong and Moskalenko, 1998; Strong et al., 2000). For -decay and bremsstrahlung, specific emissivities, using ISM H and ISM He as target, are produced which are later multiplied with the total proton distribution to obtain the final -ray emissivities in the Galaxy. For these simulations, the spatial grid is the same as in the transport step and CR energies range from 10 GeV to 1 TeV in 10 logarithmically equidistant grid points. After the diffuse emission is computed, the specific emissivities for -decay and bremsstrahlung are interpolated onto a Cartesian grid with the same spatial resolution as the sample maps, to multiply them with the total gas distribution. Such grid preserves the full resolution of the sample maps, spanning the same dimensions as that of the 3D model, with (640,640,128) grid points in the () dimensions. From the local emission, all-sky -ray emission follows via line-of-sight (LOS) integration.